Josh has 5 cousins who were born at intervals of 3 and a half years between each one.
The oldest is now exactly 3 times as old as the youngest.
How old is the middle one of the 5?

Answer by addingup(3677)

Josh has 5 cousins:
x, x+3.5, x+7, x+10.5, x+14
And
x+14 = 3x
14 = 2x
x = 14/2 = 7
And the middle cousin is x+7, so:
7+7 = 14 the middle cousin is 14 years old.